# VINP: Variational Bayesian Inference with Neural Speech Prior for Joint ASR-Effective Speech Dereverberation and Blind RIR Identification

### v2

@@ Abstract
Reverberant speech, denoting the speech signal degraded by the process of reverberation, contains crucial knowledge of both anechoic source speech and room impulse response (RIR). This work proposes a variational Bayesian inference (VBI) framework with neural speech prior (VINP) for joint speech dereverberation and blind RIR identification. In VINP, a probabilistic signal model is constructed in the time-frequency (T-F) domain based on convolution transfer function (CTF) approximation. For the first time, we propose using an arbitrary discriminative dereverberation deep neural network (DNN) to predict the prior distribution of anechoic speech within a probabilistic model. By integrating both reverberant speech and the anechoic speech prior, VINP yields the maximum a posteriori (MAP) and maximum likelihood (ML) estimations of the anechoic speech spectrum and CTF filter, respectively. After simple transformations, the waveforms of anechoic speech and RIR are estimated. Moreover, VINP is effective for automatic speech recognition (ASR) systems, which sets it apart from most deep learning (DL)-based single-channel dereverberation approaches. Experiments on single-channel speech dereverberation demonstrate that VINP reaches an advanced level in most metrics related to human perception and displays unquestionable state-of-the-art (SOTA) performance in ASR-related metrics. For blind RIR identification, experiments indicate that VINP attains the SOTA level in blind estimation of reverberation time at 60 dB (RT60) and direct-to-reverberation ratio (DRR). Codes and audio samples are available online111https://github.com/Audio-WestlakeU/VINP.
@@
Reverberation, which is formed by the superposition of multiple reflections, scattering, and attenuation of sound waves in a closed space, is one of the main factors degrading speech quality and intelligibility in daily life. The reverberant speech contains the crucial knowledge of both anechoic source speech and room impulse response (RIR). The anechoic speech, often regarded as an oracle, serves as the ultimate target of speech dereverberation. Moreover, RIR characterizes the sound propagation from sound source to microphone constrained in a room environment. Therefore, given a reverberant microphone recording, there is a strong need to estimate the anechoic speech and RIR, leading to two distinct tasks: speech dereverberation and blind RIR identification, respectively.
@@
A series of classical speech dereverberation approaches build deterministic signal models of anechoic source speech, RIR, and reverberant microphone recording. After that, dereverberation is solved by designing an inverse filter in the time domain [1] or time-frequency (T-F) domain [1, 2, 3, 4, 5]. In contrast, another series of classical methods regard the anechoic source speech signal and the reverberant microphone recording as random variables, and use hierarchical probabilistic graphical models to describe the conditional independent relationship between these random variables. Speech dereverberation is then performed by estimating every unknown latent variable and model parameter, including the anechoic speech spectral coefficient and the reverberation model parameters [6, 7]. Actually, the construction of the probabilistic graphical model is highly flexible. By modeling different factors that affect the observed signal as latent variables and model parameters, similar methods can be applied to a variety of fields, such as speech denoising [6, 8, 9] and direct-of-arrival estimation [10, 11, 12]. Classical methods are model-based and unsupervised, which makes them free from generalization problems. However, due to the insufficient utilization of prior knowledge regarding speech and distortion, the performance of these methods remains unsatisfactory.
@@
In recent decades, data-driven deep learning (DL)-based approaches have developed rapidly and surpassed the performance of classical ones. These approaches rely less on the assumptions of signal models, instead directly learn the characteristics of speech signals using deep neural networks (DNNs). Regarding such methods, the focal point of algorithm development is the design of DNN structures, network inputs and outputs, and loss functions. The most straightforward DL-based idea is to construct a discriminative DNN to learn the mapping from degraded speech to target speech. For instance, authors in [13, 14, 15, 16, 17, 18] developed various discriminative DNNs and loss functions to build mappings in various representation domains. Another concept is to consider speech dereverberation as a generative task and utilize generative DNNs, such as variational autoencoder (VAE) [19], generative adversarial network (GAN) [20], and diffusion model (DM) [21], to generate audio samples [22, 23, 24, 25]. Thanks to the powerful non-linear modeling ability of DNNs, DL-based methods are able to make the best of a large amount of data and typically lead to a better performance than classical ones. However, it is still challenging to improve automatic speech recognition (ASR) performance because such a single-channel front-end system without joint training always introduces artifact errors into speech waveforms [26, 27].
@@
Particularly, some methods combine the VAE and classical methods based on probabilistic graphical models, forming a so-called semi-supervised [28, 29] or unsupervised [30, 31] branch. Unlike the supervised DL-based approaches, the VAE in these methods is trained with only clean speech utterances. At the inference stage, the VAE decoder participates in solving the probabilistic graphical model through the estimation of clean speech prior. Usually, the Markov chain Monte Carlo (MCMC) algorithm is used to sample the latent variables in VAE. Compared with classical methods, the prior generated by VAE has higher quality and can yield better speech enhancement results. Such algorithms were first applied to speech denoising [28, 29, 30] and then expanded to speech dereverberation [31, 32]. For instance, in [31], the authors developed a Monte Carlo expectation-maximization (EM) algorithm based on a convolutional VAE and non-negative matrix factorization (NMF) model. In our previous work RVAE-EM [32], we employed a more powerful recurrent VAE to learn the prior distribution of anechoic speech. Also, we found that when the anechoic speech prior is of sufficient quality, MCMC is unnecessary, thus avoiding the repeated inference by VAE decoders. Moreover, by introducing supervised data into the training of VAE, we also obtained a supervised version of RVAE-EM that performs better than the unsupervised one. This finding inspired our current work. Since there are already many advanced dereverberation DNN architectures, it is possible to directly apply them as supervised anechoic speech prior estimators to the solution of the probabilistic graphical model with some simple modifications.
@@
Within the domain of audio signal processing, blind RIR identification from reverberant microphone recording constitutes a crucial and challenging area of research, since RIR characterizes the acoustic characteristics of the sound propagation and environment. Currently, methods for blind RIR identification are rather scarce. In recent years, with the development of DL techniques, some DL-based approaches have been proposed [33, 34, 35]. For instance, the authors in [35] established a parameterized model for the reverberation effect and proposed an unsupervised method for joint speech dereverberation and blind RIR estimation. This approach is similar to our work in terms of tasks.
@@
In this paper, we propose a variational Bayesian inference framework with neural speech prior (VINP) for joint speech dereverberatiion and blind RIR identification. Our motivation is as follows: Given that the convolution transfer function (CTF) approximation is a rather accurate model of the reverberation effect, by making reasonable probabilistic assumptions and regarding the output of DNN as the prior distribution of anechoic speech, we are able to analytically estimate the anechoic spectrum and the CTF filter, and further obtain the time-domain waveforms of anechoic speech and RIR. The basis of VINP is a probabilistic signal model based on CTF approximation [36], the same as in our previous work RVAE-EM [32]. The signal model describes the relationship among the source speech, CTF filter, and reverberant microphone recording. Different from RVAE-EM [32], considering the existence of advanced discriminative DNN structures for dereverberation, we proposed to employ such DNNs to learn the prior distribution of anechoic speech by modifying the loss function of network training. Sequentially, we use variational Bayesian inference (VBI) [37, 38] to analytically solve the hidden variables and parameters in our probabilistic model. By doing this, VINP avoids the direct utilization of DNN output but still utilizes the powerful nonlinear modeling capability of the network, thereby benefiting ASR. Moreover, a major drawback of RVAE-EM [32] is that the computational cost grows with the cube of the speech length. Through the variational inference, the computational cost in the proposed method increases linearly with the speech length. Additionally, parallel computation can be used across T-F bins for fast implementation.
@@
We propose VINP, a novel VBI framework with neural speech prior for joint speech dereverberation and blind RIR identification. For the first time, we propose introducing an arbitrary discriminative dereverberation DNN into VBI to successfully complete the tasks.
VINP avoids the direct utilization of DNN output but still utilizes the powerful nonlinear modeling capability of the network. Thus, it can enhance the ASR performance of the original backbone DNN without the need for any joint training with the ASR system. Experiments demonstrate that VINP attains an advanced level in most metrics related to human perception and displays unquestionable state-of-the-art (SOTA) performance in ASR-related metrics.
VINP can be used for blind RIR identification from reverberant microphone recording. Experiments indicate that VINP attains the SOTA level in the blind estimation of reverberation time at 60dB (RT60) and direct-to-reverberation ratio (DRR).
From the perspective of computational cost, VINP exhibits linear growth with respect to the speech length, which is different from the cubical growth in our previous work RVAE-EM [32]. Moreover, the VBI procedure in VINP can be implemented in parallel across T-F bins, thereby further reducing the processing time.
II Signal Model and Task Description
In this section, we will introduce the signal model and define two tasks we aim to address: speech dereverberation and blind RIR identification.
@@ II-A Signal Model
Considering the scenario of a single static speaker and stationary noise, the reverberant speech signal (observation) received by a distant microphone can be modeled in the time domain as
where is the convolution operator, is the index of sampling points, is the observation signal, is the anechoic source speech signal, is the RIR which describes a time-invariant linear filter, and is the background additive noise. Without loss of generality, we assume that the RIR begins with the impulse response of the direct-path propagation, followed with reflections and reverberation.
Analyzing and processing speech signals in the time domain poses significant challenges. After performing short-time Fourier transform (STFT), according to the CTF approximation [36], the observation model in the T-F domain becomes
@@
Furthermore, the observation , anechoic source , and noise are modeled as random signals. We have the following assumptions regarding their distributions.
Assumption 1: The anechoic source speech signal follows a time-variant zero-mean complex-valued Gaussian distribution, while the noise signal follows a time-invariant zero-mean complex-valued Gaussian distribution. Therefore, we have their prior distributions as
@@
Assumption 2: The anechoic source speech signal and noise signal are respectively independent for all T-F bins. Defining and , we have
@@
Assumption 3: The anechoic source speech signal and noise signal are independent of each other, which means
With all the aforementioned assumptions, the conditional probability of the observation signal can be expressed as follows.
The probabilistic graphical model is summarized in Fig. 1.
@@ II-B Task Description
In this work, we aim to jointly complete speech dereverberation and blind RIR identification by solving all hidden variables and parameters in Fig. 1 through VBI.
For speech dereverberation, we care about the anechoic spectrum, which is a hidden variable in our probabilistic graphical model. We aim to get its maximum a posteriori (MAP) estimation given the reverberant microphone recording, written as
where the posterior distribution of anechoic source signal can be represented according to the Bayes rule as
For blind RIR identification, we care about the CTF filter, which is a model parameter in our probabilistic graphical model. Defining the CTF filter for all frequency bands as , we aim to get its maximum likelihood (ML) estimation, written as
The CTF filter is a representation of RIR in the T-F domain. We transform the CTF filter into the RIR waveform through a pseudo measurement process, a process that will be elaborated upon later.
@@ III Proposed Method
In this work, we propose a novel framework named VINP for joint speech dereverberation and blind RIR identification. We propose using an arbitrary discriminative dereverberation DNN to predict the prior distribution of anechoic speech from reverberant microphone recording and then applying VBI to analytically estimate the anechoic spectrum and the CTF filter. After that, we use a pseudo measurement process to transform the CTF filter into the RIR waveform. The overview of VINP is shown in Fig. 2.
III-A Prediction of Anechoic Speech Prior
The direct-path speech signal, which is a scaled and delayed version of the anechoic source speech signal, is free from noise and reverberation as well. To avoid estimating the arbitrary direct-path propagation delay and attenuation, instead of the actual source speech, we setup the direct-path speech as the source speech and our target signal, still denoted as or . Correspondingly, the RIR begins with the impulse response of the direct-path propagation.
In VINP, we consider the power spectrum of the direct-path speech signal as the variance of the anechoic speech prior . Then, in each T-F bin, the oracle estimation of precision is
However, the oracle estimation is unavailable in practice. We propose integrating an arbitrary discriminative dereverberation DNN as an estimator of anechoic speech prior. Note that, in the proposed framework, we need to redesign the training loss function of the discriminative dereverberation DNNs.
Specifically, the discriminative DNN constructs a mapping from reverberant magnitude spectrum to the anechoic magnitude spectrum as
Then Eq. (10) becomes
Regarding the loss function, we employ the average Kullback-Leibler (KL) divergence [39] to measure the distance of estimated prior distribution and oracle prior distribution as
In practice, we use
instead to avoid numerical instabilities, where is a small constant. Such a loss function is different from that in regular discriminative dereverberation DNNs which directly predict the anechoic spectrum.
Research indicates that the complex nonlinear operations of DNNs often lead to outputs with unpredictable artificial errors. While these errors do not significantly affect speech perceptual quality, their impact on back-end speech recognition applications remains uncertain. As a result, DL-based approaches may not improve ASR performance [27, 40, 26]. However, in our method, by regarding the DNN output as a prior distribution of anechoic speech and utilizing the subsequent CTF-based VBI stage to refine it, this problem can be alleviated, leading to better ASR performance. Experiments in Section IV will provide evidence to support this conclusion.
@@ III-B Variational Bayesian Inference
Given the estimated prior distribution of anechoic speech and the observed recording, the estimation of every hidden variable and parameter is carried out through variational Bayesian inference.
The posterior distribution is intractable due to the integral term in Eq. (8). Therefore, we turn to VBI, which is a powerful tool for resolving hierarchical probabilistic models [11, 41]. More specifically, we employ the variational expectation-maximization (VEM) algorithm, which provides a way for approximating the complex posterior distribution with a factored distribution according to the mean-field theory (MFT) as
After factorization, VEM can estimate the posterior and model parameters by E-step and M-step respectively and iteratively as
@@
where denotes the set subtraction and denotes expectation. Because the solution of such a probabilistic graphical model is an underdetermined problem, we do not update the precision parameter during the VEM iterations to prevent a deterioration in prior quality.
The specific update formulae are as follows:

### v3

@@ Abstract
[…]
@@
Reverberation, which is formed by the superposition of multiple reflections, scattering, and attenuation of sound waves in a closed space, is one of the main factors degrading speech quality and intelligibility in daily life. The reverberant speech contains the crucial knowledge of both anechoic source speech and room impulse response (RIR). The anechoic source speech is clean and serves as the target of the dereverberation task (with possible time delay and scaling). Moreover, RIR characterizes the sound propagation in an enclosure. An accurate estimate of RIRs contributes to applications including speech recognition [1], speech enhancement [2] and augmented/virtual reality [3]. Therefore, given a reverberant microphone recording, there is a strong need to estimate the anechoic speech and RIR, leading to two distinct tasks: speech dereverberation and blind RIR identification, respectively.
@@
A series of classical speech dereverberation approaches build deterministic signal models of anechoic source speech, RIR, and reverberant microphone recording. After that, the task is solved by designing an inverse filter in the time domain [4] or time-frequency (T-F) domain [4, 5, 6, 7, 8]. In contrast, another series of classical methods regard the anechoic source speech and the reverberant microphone recording as random signals, and use hierarchical probabilistic models to describe the generation process of the observation. After that, speech dereverberation is performed by estimating every unknown part in the probabilistic model, including hidden variables and model parameters [9, 10]. The construction of the probabilistic model is highly flexible. Similar models can be applied to a variety of fields, such as speech denoising [9, 11, 12] and direct-of-arrival estimation [13, 14, 15]. These classical approaches are model-based, which makes them free from generalization problems. However, due to the insufficient utilization of prior knowledge regarding speech and reverberation, their performance remains unsatisfactory.
@@
In recent decades, data-driven deep learning (DL)-based approaches have developed rapidly. These approaches rely less on the assumptions of signal models, instead directly learn the characteristics of speech signals using deep neural networks (DNNs). The core research of DL-based approaches focus in the design of DNN structures, features, and loss functions. The most straightforward DL-based idea is to construct a discriminative DNN to learn the mapping from degraded speech to target speech. For instance, authors in [16, 17, 18, 19, 20, 21] developed various discriminative DNNs and loss functions to build mappings in various feature domains. Another idea is to consider speech dereverberation as a generative task and utilize generative DNNs, such as variational autoencoder (VAE) [22], generative adversarial network (GAN) [23], and diffusion model (DM) [24], to generate anechoic speech [25, 26, 27, 28]. Thanks to the powerful non-linear modeling ability of DNNs, DL-based methods are able to make the best of a large amount of data and typically lead to better perceptual performance than classical approaches. However, when applied as single-channel front-end systems without joint training, these DL-based methods may introduce waveform artifacts that degrade subsequent automatic speech recognition (ASR) performance [29, 30].
@@
Particularly, some approaches combine DNNs and probabilistic signal models, forming a so-called semi-supervised [31, 32] or unsupervised [33, 34] paradigm. Unlike fully supervised DL-based approaches, these methods train VAEs with only clean speech corpora. At the inference stage, the VAE decoder participates in solving the probabilistic model by estimating the prior distribution of clean speech. Usually, the Markov chain Monte Carlo (MCMC) algorithm is used to sample the latent variables in VAE. Compared with classical methods, the prior generated by VAE has higher quality and can yield better performance. Such algorithms are applied to both speech denoising [31, 32, 33] and speech dereverberation [34, 35]. For instance, in [34], the authors developed a Monte Carlo expectation-maximization (EM) dereverberation algorithm based on a convolutional VAE and non-negative matrix factorization (NMF). In our previous work RVAE-EM [35], we employed a more powerful recurrent VAE and found that when the anechoic speech prior is of sufficient quality, MCMC is unnecessary, thus avoiding the repeated inference of VAE decoders. Moreover, by introducing supervised data into the training of VAE, we also obtained a supervised version of RVAE-EM that performs better than the unsupervised one. This finding inspired our current work. Given the availability of many advanced dereverberation DNNs, it is possible to directly apply them as supervised prior estimators in the probabilistic model solution through straightforward modifications.
@@
Within the domain of audio signal processing, blind RIR identification from reverberant microphone recordings is a crucial and challenging area of research. Currently, methods for blind RIR identification are rather scarce. In recent years, with the development of DL techniques, some DL-based approaches have been proposed [36, 37, 38]. For instance, in FiNS [36], the authors utilized DNN to learn the direct path, early reflection, and late reverberation of RIR separately. The authors in BUDDy [38] established a parameterized model for the reverberation effect and proposed an unsupervised method for joint speech dereverberation and blind RIR estimation.
@@
In this paper, we propose a Variational Bayesian Inference framework with Neural speech Prior (VINP) for joint speech dereverberation and blind RIR identification. Our motivation is as follows: The generation of reverberant recording can be modeled in a probabilistic model based on convolution transfer function (CTF) approximation [39, 40]. By treating the DNN output as the prior distribution of anechoic speech and combining it with the reverberant recording to solve the probabilistic signal model, we can analytically estimate the anechoic spectrum and the CTF filter, and subsequently obtain the anechoic speech and RIR waveforms. By doing this, VINP avoids the direct utilization of DNN output but still leverages the powerful nonlinear modeling capability of DNN, and further ensures that the estimates and observations align with the CTF-based signal model, thereby benefiting ASR. Different from our previous work RVAE-EM [35], in VINP, we propose to employ an arbitrary discriminative DNN to learn the prior distribution of anechoic speech by modifying the loss function during training. Moreover, a major drawback of RVAE-EM is that the computational complexity scales cubically with speech duration. In VINP, we use variational Bayesian inference (VBI) [41, 42] to analytically solve the probabilistic model. Thanks to VBI, the computational complexity in VINP increases linearly with the speech duration. Parallel computation is available across T-F bins for efficient implementation.
@@
We propose VINP, a novel framework for joint speech dereverberation and blind RIR identification. For the first time, we propose introducing an arbitrary discriminative dereverberation DNN as a backbone into VBI to successfully complete these two tasks.
VINP avoids the direct utilization of DNN output but still utilizes the powerful nonlinear modeling capability of the network, and further ensures that the estimates and observations align with the CTF-based signal model. As a result, VINP attains state-of-the-art (SOTA) performance in mean opinion score (MOS) and word error rate (WER).
VINP can be used for blind RIR identification. Experiments demonstrate that VINP attains SOTA and advanced levels in the blind estimation of reverberation time at 60dB (RT60) and direct-to-reverberation ratio (DRR).
From the perspective of computational complexity, VINP exhibits linear scaling with respect to the speech duration, representing a significant improvement over the cubical scaling in our previous work RVAE-EM [35]. Moreover, the VBI procedure enables parallel processing across T-F bins, which further accelerates inference.
II Signal Model and Tasks
@@ II-A Signal Model
Considering the scenario of a single static speaker and stationary noise, the reverberant recording (also known as observation) received by a distant microphone can be modeled in the time domain as
where is the convolution operator, is the index of sampling points, is the observation signal, is the anechoic source speech signal, is the RIR which describes a time-invariant linear filter, and is the background additive noise.
Analyzing and processing speech signals and the very long RIR filter (normally thousands of taps) in the time domain poses significant challenges. For example, the estimation of filter, e.g. Eq. (25) in the proposed method, shares similar solutions in T-F domain and time domain. But the solution in time domain requires the inversion of a very large correlation matrix, which suffers from very high computational/memory costs and poor robustness. Therefore, performing short-time Fourier transform (STFT), we analyze the signals in T-F domain, in which reverberation can be modeled using a series of cross-band filters [39]. Since the energy of cross-band filters is concentrated and more crossband filters may lead to a higher variance in estimation error [39], we consider only band-to-band filters to simplify the analysis, resulting in CTF approximation [40]. Theoretical and experimental error analysis of CTF approximation appears in [39] and [43]. The observation in the T-F domain is
@@
Furthermore, the observation , the anechoic source , and noise are modeled as random signals. We have the following assumptions regarding their prior distributions.
Assumption 1: The anechoic source speech follows a time-variant zero-mean complex-valued Gaussian prior distribution, while the noise signal follows a time-invariant zero-mean complex-valued Gaussian prior distribution. Therefore, we have
@@
Assumption 2: Both the anechoic source speech signal and the noise signal are statistically independent across all T-F bins. Defining and , we have
@@
Assumption 3: The anechoic source speech and noise are independent of each other, which means
With all the aforementioned assumptions, the conditional probability of the observation signal can be expressed as
where and . The probabilistic graphical model, which describes the generation process of reverberant recording, is shown in Fig. 1.
@@ II-B Task Description
Our goal is to jointly achieve speech dereverberation and blind RIR identification by solving all hidden variables and parameters in Fig. 1 via VBI.
For speech dereverberation, our focus is on the anechoic speech spectrum, which corresponds to the hidden variables in our probabilistic graphical model. We aim to get its maximum a posteriori (MAP) estimate given the reverberant recording, written as
where the posterior distribution of anechoic source signal can be represented according to Bayes’ rule as
For blind RIR identification, our focus is on the CTF filter, which corresponds to the model parameters in our probabilistic model. Defining the CTF filter in all frequency bands as , we aim to obtain its maximum likelihood (ML) estimate, written as
The CTF filter provides a representation of RIR in the T-F domain. We reconstruct the RIR waveform from CTF coefficients through a pseudo measurement process, as detailed in Section III.
@@ III Proposed Method
In VINP, we propose using an arbitrary discriminative dereverberation DNN to estimate the prior distribution of anechoic speech and using VBI to analytically estimate the anechoic spectrum and CTF filter. Subsequently, we use a pseudo measurement process to transform the CTF filter into the RIR waveform. The overview of VINP is shown in Fig. 2.
III-A Estimation of Anechoic Speech Prior
The direct-path speech signal, which is a scaled and delayed version of the anechoic source speech signal, is free from noise and reverberation. To avoid estimating the arbitrary direct-path propagation delay and attenuation, instead of the actual source speech, we setup the direct-path speech as the source speech and our target signal, which is still denoted as or . Correspondingly, the RIR begins with the impulse response of the direct-path propagation.
In VINP, we consider the periodogram of the direct-path speech signal as the variance of the anechoic speech prior . In each T-F bin, the ideal precision parameter in Eq. (3) is
Because Eq. (10) is unavailable in practice, we propose using an arbitrary discriminative dereverberation DNN to estimate the anechoic speech prior . DNN constructs a mapping from reverberant magnitude spectrum to the anechoic magnitude spectrum as
Correspondingly, the estimated anechoic speech prior becomes
where
Similar to , is an element in that corresponds to frequency band and frame .
Regarding the loss function, we employ the average Kullback-Leibler (KL) divergence [44] to measure the distance of the estimated prior distribution and oracle prior distribution as
where is a small constant to avoid numerical instabilities. Such a loss function differs from that in conventional discriminative dereverberation DNNs, which directly estimate the anechoic spectrum, yet it shares the same formulation as the Itakura-Saito divergence between spectrograms employed in [33, 34, 35, 45]. In preliminary experiments, we have also explored various loss functions, including mean squared error (MSE) and mean absolute error (MAE) for linear and logarithmic magnitude spectra, alongside backward KL divergence. The experimental results confirm that the KL loss outperforms all the alternatives.
Following the DNN, approximations and are employed in the subsequent VBI procedure. Notice that the independent assumption of anechoic speech prior is merely intended to simplify the subsequent derivation of VBI. The DNN learning of speech prior is not limited to the exclusion of correlations between T-F bins.
@@ III-B Variational Bayesian Inference
Given the estimated prior distribution of anechoic speech and the observation, we solve all hidden variables and parameters via Bayesian inference.
The posterior distribution is intractable due to the integral term in Eq. (8). Therefore, we turn to VBI, a powerful tool for resolving hierarchical probabilistic models [46]. More specifically, we employ the variational expectation-maximization (VEM) algorithm, which provides a way for approximating the complex posterior distribution with a factored distribution according to the mean-field theory (MFT) as
After factorization, VEM estimates the posterior and model parameters by E-step and M-step respectively and iteratively as
@@
where denotes the set subtraction and denotes expectation. For simplicity, in the subsequent descriptions, we omit the subscript of . We do not update the precision parameter during iterations to prevent the degradation of prior distribution.
The specific formulas are as follows: